Amber sells athletic shoes part-time at a department store. She can earn either $500 per month plus a 8% commission on her total sales, or $400 per month plus a 12% commission. What is the total price of the athletic shoes that she needs to sell to earn the same income from each pay scale?
Answer by rfer(16322)

0.08x+500=0.12x+400
0.08x-0.12x=400-500
-0.04x=-100
x=-100/-0.04
x=2500